# Zero-Shot Machine-Generated Text Detection Using Mixture of Large Language Models

### v1

@@ 3.2 Baselines
Machine-generated text detection methods are usually divided into two main categories (Dugan et al., 2024): supervised and unsupervised (metric-based). The former uses supervision data to fine-tune a pre-trained model for the detection task, typically focusing on some known generators. Our method belongs to the latter family, as our metric (the RSA score of 2) does not require any training data.
Accordingly, we compare our method to other zero-shot unsupervised techniques used for machine-generated text detection:
@@
  • Binoculars score (Hans et al., 2024) is a perplexity-based method detailed in Eq. (23) in the Appendix. With this score, generated texts should score lower than human-produced ones. We use the default parameters given by the authors, chosen for their best results on the Binoculars dataset, with falcon-7b-instruct as the detector model, used to calculate perplexity, and falcon-7b as the auxiliary model.
  • DetectGPT (Mitchell et al., 2023) generates minor perturbations of a given text passage, then calculates the differences in log-probabilities between the original text and its perturbed versions. Looking at the curvature of the log-probability function, text from LLMs show less variability and remains in regions of negative curvature when slightly perturbed compared to human-generated text. We used the default parameters provided, with GPT2-XL computing the log probabilities and T5-3b generating the perturbations. To better fit the implementation, we truncated our texts when necessary and selected a random sample of 1,000 pairs of original and generated texts. We run the "100_d" experiment with 100 perturbations applied to each text.
  • Measuring the perplexity (PPL). The text’s log-perplexity as seen by all models in our ensemble –as LLMs usually generate text having a lower perplexity than human’s productions (Vasilatos et al., 2023, Guo et al., 2023, Mitrović et al., 2023, Li et al., 2024)– leads to a very straightforward metric for detecting machine-generated text. PPL (oracle) means we chose the best discriminating (codelength) model, PPL (average) is how good the score is at discriminating with the average of all perplexities.
All the scores reported in this study are ours, using the implementation provided in the original papers.
4 Numerical Results
@@
The following numerical results all use the same configuration, RSA-4, unless explicitly stated otherwise. Our default ensemble of models is composed of TowerBase-7b, TowerBase-13b, Llama-2-7b-chat and Llama-2-7b. All AUC scores are obtained using scikit-learn (Pedregosa et al., 2011).
4.1 Results for the Binocular Benchmark
4.1.1 Baseline ensemble
Our main results are in Table 2, where we report detection scores for the Binoculars datasets (both the original version and our Llama-base addition (results with Mistral-based generations are in Appendix C). For the first, the best two approaches are Binoculars and RSA which outperform the other baselines. Note that contrarily to Binoculars, our results are obtained without any tuning, and without using the generator model in our ensemble. RSA-4 (with averaging, our best variant here) even slightly outperforms this baseline on Pubmed, whose vocabulary is more specialized towards medicine than the one used in CNN and CC News. Results on CNN are harder to interpret, as artificial texts in this dataset appear to be easily uncovered: using a threshold on the perplexity of our models leads to a near-perfect identification. In the following, we ignore this sub-dataset as all methods perform equally well on it.
Pubmed CNN CC_news
Original Regen Original Regen Original Regen
PPL (oracle) 0.823 0.690 0.598 1.000 0.617 0.632
PPL (average) 0.808 0.637 0.566 0.999 0.584 0.569
Detect GPT 0.716 0.668 0.642 0.458 0.580 0.570
Binocular score 0.988 0.633 0.995 1.000 0.979 0.578
RSA-4 (avg) 0.992 0.887 0.993 0.999 0.971 0.854
RSA-4 (max) 0.979 0.904 0.992 0.993 0.948 0.861
RSA-5 (+phi) 0.992 0.885 0.988 0.998 0.950 0.815
Table 2: Detecting artificial texts in the Binoculars Dataset (AUC scores).
4.1.2 Including the generator in our ensemble
For the Llama-7b-based regeneration, we see a large decrease of Binoculars’s performance, as its results become comparable to the PPL thresholding baselines. We note that the (max) variant of robust scoring algorithm becomes the best-performing approach, presumably owing to the strong predictions of the Llama-2 models. If RSA’s results are more stable than Binoculars, the decrease is still significant for Pubmed and CC_news.
This decrease may seem surprising since Llama-2-7b is one of the models we use to detect the generated text. We hypothesise that the sampling method plays here a very important role in the performance of our methods. The Binoculars dataset was generated using a temperature of 0.70.70.70.7777Personal correspondence with the first author., while our regeneration used a temperature of 0.6 along with nucleus sampling ($p=0.9$) and a repetition penalty of 1.18. These parameters are recommended by the official Llama repository [^8] https://github.com/meta-llama/llama/blob/main/example_text_completion.py as they should produce results more diverse, and in a way, human-like than the logits output by raw models. This means that the actual sampling distribution does not fully match the base distribution used in the detector.
In a follow-up experiment, we replace the distribution computed by Llama-7b with a “distorted” version, which approximates the effects of temperature and nucleus-sampling [^9] For this, we first divide the model logits by the temperature, then apply the softmax and perform nucleus filtering with $p=0.9$, making sure to smooth the resulting distribution so that filtered tokens still get a small positive value.. This has a clear effect on the Blahut–Arimoto weights ($q^{*})$, which increase for Llama-7b and decrease for Llama-7b-instruct, but hardly change the detection performance (see Figure 1 in Appendix). This hints at the fact that for generator-detector approaches, the actual generator sampling policy should also be taken into account.
4.1.3 Augmenting the ensemble
RSA makes the augmentation of the ensemble quite easy, as long as all models use the same underlying tokenizer. To showcase the effect of this feature, we add Phi-3 (Abdin et al., 2024) in our ensemble. In this experimental setting, this extension of the model set is of little consequence (see line “+phi” in Table 2), the most notable variation being a small drop in performance for the regenerated CC_news subpart (4% of AUC in the worst-case scenario). Yet, we observe that Phi consistently gets a substantial share of the Blahut–Arimoto weights (see Figure 2(b) in Appendix) and plays a significant role in the classification decision, which has the effect of making Llama-generated texts slightly less likely for the combined model.
4.1.4 Handling poor generators: the case of unigram models
As noted by Hans et al. (2024), simple random generators are hard to detect for Binoculars (as well as for detectors using a PPL-based threshold): this is because random word salads are “surprising” (have a high perplexity) for well-trained detectors, and tend to be confused with human productions. We reproduce this observation using the Random corpus and observe that all baselines, as well as our proposed detector, achieve AUC scores close to 00 for this dataset. The PPL of word salads is however much larger than for human texts, so setting an upper bound of the human PPL would, in that case, provide a very easy fix.
It is tempting to see whether adding such a poor generator into our combination is of any help. For this experiment, we train a unigram language model [^10] We process the Brown Corpus (Francis and Kucera, 1979), tokenized with Llama tools and use count ratios to estimate the probabilities of each token. A smoothing parameter of $\epsilon=1e-10$ is added to each count to avoid null values. and combine it with our four baseline detectors.
The detector obtained with this extended ensemble remains completely unable to sort random from human texts: having the random model in the robust scoring algorithm does not make random texts more likely. This model is also slightly detrimental when it comes to detecting strong generators, as we observe a mild drop in performance throughout (see Appendix, Table 4 for the Binoculars dataset). This is because the unigram model predicts unexpected tokens most of the time and therefore often gets a significant weight in the Blahut–Arimoto weights (see Figure 2(c) in Appendix), which leads to a huge discrepancy between the cross-entropies scores that are averaged in Definition 1.
4.2 Detecting ChatGPT: Ghostbuster and M4
For Ghostbuster, a threshold on the perplexity is enough to achieve an AUC over $0.99$. As a matter of fact, those ChatGPT generations are quite easy to detect. Full results for this dataset are given in Table 7 (Appendix D), where we see that Binoculars gets the highest scores, while RSA-4 and PPL based baselines also give very good AUC scores.
For multilingual data, however, the task is much harder. We considered for this experiment the Arabic, Russian, and Bulgarian datasets from the M4 paper Wang et al. (2024a), as these languages are better represented in the training set of the Tower Family of LLMs than in the Llama-2 family (used in our ensemble) and the Falcon family (used by the Binoculars score). Having multilingual detector models enables us to obtain good performance for Arabic, Bulgarian, and Chinese, less so for Russian. Binoculars scores remain strong for Chinese, but tend to drop for the other languages. When it comes to model selection, our method appears to give more weight to the TowerBase models in our ensemble compared to when just looking at English texts (see Figure 4 in Appendix). This nicely illustrates the benefits of using multiple and complementary models, each one with a restricted domain of “expertise”.
Method Arabic Russian Bulgarian
PPL (oracle) 0.918 0.694 0.932
BinocularsScore 0.686 0.674 0.742
RSA (ours) 0.909 0.737 0.974
Table 3: AUC for Multilingual ChatGPT.
4.3 Detecting scholarly texts
@@
As previously mentioned, our dataset of scholarly texts is composed of fully generated papers. As they were produced by a GPT-2 model fine-tuned on Arxiv’s “computation and language” section, the distribution of tokens greatly differs from all models considered in this study. Therefore, the generated texts are completely out-of-domain and get a high perplexity, which makes them difficult to discriminate from human texts (see Figure 3 in the Appendix) with perplexity-based models. This holds, in particular, for RSA-4 and Binoculars, which achieve respective AUCs of 0.421 and 0.505. This shows that brittleness issues are not fully solved, and hints at including more domain-adapted models in detector ensembles.
@@ 5 Related Work
The improvement of text generation abilities of LLMs has raised many concerns about potential misuses such as generating disinformation (Zellers et al., 2019), abusive content (Crothers et al., 2023), forged academic publications (Liu et al., 2024), or to cheat during academic exams (Vasilatos et al., 2023). Such fake texts seem difficult for humans to spot (Gehrmann et al., 2019). In response, the issue of automatically detecting machine-generated texts has been subject to an increasing focus. This problem can be framed in multiple ways: as a binary (human vs. non-human) decision problem, as the problem of detecting one known artificial agent (e.g.,  ChatGPT (Mitrović et al., 2023; Liu et al., 2024)), or as discriminating the correct model in a predefined list (Li et al., 2023). As texts are extremely varied, another distinction is between closed-domain (e.g. scientific (Liyanage et al., 2022), academic (Liu et al., 2024) or user-generated content (Fagni et al., 2021; Kumarage et al., 2023)) vs. open-domain text detection. Assuming the generator models are known, various settings can be considered, depending on whether models can be openly queried (open parameters), whether they expose their full logits, or just the top prediction (and associated probability), etc.
@@
Supervised detection, especially when one single generator is used, looks like a simple problem, as detection rates in the high 90s are easily achieved (Zellers et al., 2019; Guo et al., 2023; Liu et al., 2024) using strong text classifiers based on features extracted by Roberta (Conneau et al., 2020) or T5 (Raffel et al., 2020). These approaches are brittle, though, and their success depends on particular generator-detector pairs (Antoun et al., 2024), prompting e.g. Verma et al. (2023) to design automatic feature extractors from an ensemble of detectors to improve the robustness of their system.
@@
Unsupervised detection is somewhat more challenging. Most approaches to date rest on the intuition that human-written texts are more “surprising” than machine-generated texts [^11] Assuming generation does not use random sampling, in which case the reverse is likely to be observed, as long artificial texts drift away from natural writings (Zellers et al., 2019)., which should translate into a tangible difference in token-wise log-probability121212(Mitchell et al., 2023) argues that the difference is better seen at the level of log-ranks. between the two. This idea is already used in GPTzero [^13] https://gptzero.me/ and thresholding perplexity usually provides strong baselines (see, inter alia, (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023)). Such techniques heavily rely on the detector model(s) used to compute the log-likelihood of artificial and human-generated texts, which must be robust to variations in text domains, genres, styles, and, needless to say, languages (Wang et al., 2024a); as well as variations in the generator itself Antoun et al. (2024).
@@
Mitchell et al. (2023) and Bao et al. (2024) exploit a similar intuition, arguing that small random perturbations [^14] Performed by a strong generation model, T5 Raffel et al. (2020) in their study. of an artificial text will on average make it less likely (more surprising), which is not true for human-written texts. They derive a statistical criterion to implement this idea, based on an estimate of the curvature of the log-probability function, and report near-perfect detection scores on three types of texts, generated by 5 models. As explained above, Hans et al. (2024) also rely on a function of the per-token log-perplexity, in their case contrasted with the cross-entropy of an auxiliary model.
@@
Abandoning generator-detector-based techniques altogether, (Mao et al., 2024; Yang et al., 2024) develop effective detection approaches based on regeneration strategies, prompting the (known) generator to regenerate part of the input text. The intuition is that artificial inputs are likely to be regenerated exactly, while human texts exhibit greater redundancy, resulting in the need for longer codelengths. Other strategies to combat AI-generated text detection have also been considered. Text watermarking (Kirchenbauer et al., 2023a, b; Liu and Bu, 2024), is one of them. Its efficiency and robustness, however, are still subject to discussions, e.g., (Zhang et al., 2023).
Several recent works have focused on the robustness of detection methods. When subjecting machine-generated text to simple modifications, in (Wang et al., 2024b), only watermarking remains able to accurately identify artificial documents. Dugan et al. (2024) present a diverse set of artificial texts, generated with multiple models and sampling strategies, and additionally subject to various adversarial attacks. They also evaluate the robustness of several standard detectors with respect to these source variations, observing that most suffer large drops in performance. In their comparison, Binoculars (Hans et al., 2024) stands out, achieving decent detection scores at False Positive Rates under 1%. Some training strategies have been proposed by Wang et al. (2024b) and Huang et al. (2024) to improve on these robustness issues.

### v2

@@ 3.2 Baselines
Machine-generated text detection methods are usually divided into two main categories (Dugan et al., 2024): supervised and unsupervised (metric-based). The former uses supervision data to fine-tune a pre-trained model for the detection task, typically focusing on some known generators. Our method belongs to the latter family, as the MOSAIC score of Eq. (2), which serves to discriminate forged content, does not require any training data. Accordingly, we compare our method to other zero-shot unsupervised techniques used for machine-generated text detection:
Perplexity (PPL) based detectors use a threshold on the text’s log-perplexity, assuming that LLMs usually generate texts that have a lower perplexity than human’s productions (see Vasilatos et al., 2023; Guo et al., 2023; Mitrović et al., 2023; Li et al., 2024 inter alia). This yields a very straightforward criterion for detecting machine-generated texts. We compute this baseline separately for all available models in our ensemble. Additionally, PPL (average) reports the performance obtained with an average of all perplexities within our ensemble.
DetectGPT (Mitchell et al., 2023) generates minor perturbations of a given text passage, then computes the differences in log-probability between the original text and its perturbed versions. DetectGPT relies on the observation that when slightly perturbed, the log-probability of artificial texts consistently decreases, which is not the case for human-generated texts. We used the default parameters in the authors’ implementation, running the "10d" experiment with GPT2-medium computing the log-probabilities and T5-large (Raffel et al., 2020) generating the perturbations. Since the original implementation is not optimized for long sequences and imbalanced datasets, we split the lengthy texts into smaller chunks and downsampled the larger dataset to ensure both were of equal size. FastDetectGPT (Bao et al., 2024) relies on the same principles but uses a different approach to (quickly) sample the perturbations of the text. Its authors found that it was both faster and slightly better than the original implementation.
@@
Binoculars score (Hans et al., 2024) is also based on log-probabilities: it compares the average log-probability of an input text for a detector model with the cross-entropy of an auxiliary model (see Eq. (23) in Appendix C). With this score, artificial texts should have a score lower than natural texts. We used the default model pair selected by the authors, chosen for their best results on the Binoculars dataset, with Falcon-7b-instruct as the detector model, used to compute the perplexity, and Falcon-7b as the auxiliary model. We also report detection results using Eq. (23) with two models from our ensemble.
Baseline scores in Table 2 are obtained with the implementations provided with the original papers. These results do not directly compare with those of Table 3 as the underlying set of models is different.
Table 2: Detection performance of baseline systems. AUROC scores.
RAID Binoculars Ghostbuster M4 (multilingual) Scho. Avg.
Pubmed CNN CCnews Reuter Essay Reddit Ara Bul Ger Rus
DetectGPT 0.632 0.666 0.635 0.571 0.714 0.916 0.757 0.576 0.589 0.524 0.597 0.440 0.635
FastDetectGPT 0.706 0.787 0.925 0.772 0.829 0.949 0.938 0.874 0.683 0.827 0.596 0.549 0.786
Binoculars 0.853 0.988 0.995 0.979 0.993 0.996 0.990 0.686 0.742 0.914 0.674 0.505 0.860
Table 3: Artificial text detection performance of detectors built with a fixed set of 4 models. Detection may involve running 1, 2, or 4 models. AUROC scores.
RAID Binoculars Ghostbuster M4 (multilingual) Scho. Avg
Pubmed CNN CCnews Reddit Reuter Essay Ara Bul Ger Rus
Best single-model 0.834 0.999 0.995 0.975 0.878 0.886 0.818 0.985 0.988 0.832 0.816 0.517 0.877
Best two-model 0.803 0.989 0.994 0.973 0.677 0.663 0.481 0.897 0.959 0.860 0.492 0.381 0.764
avg PPL 0.730 0.808 0.566 0.584 0.980 0.980 0.992 0.897 0.887 0.638 0.598 0.516 0.765
$q^{\star}$ (log-probs) 0.746 0.807 0.566 0.579 0.985 0.984 0.994 0.893 0.901 0.640 0.592 0.509 0.766
MOSAIC-4 (avg) 0.850 0.992 0.993 0.971 0.946 0.971 0.911 0.909 0.974 0.890 0.737 0.421 0.880
MOSAIC-4 (unif) 0.844 0.992 0.995 0.975 0.920 0.951 0.876 0.909 0.974 0.893 0.745 0.416 0.874
4 Experimental Results
@@
The following experimental results all use the same configuration. Unless explicitly stated otherwise, MOSAIC-4 uses an ensemble of models composed of TowerBase-7b, TowerBase-13b (Alves et al., 2024), Llama-2-7b-chat, and Llama-2-7b (Touvron et al., 2023). This choice of models is motivated by their shared tokenizer.
4.1 The robustness of ensemble methods
In our experiments, we first evaluate the added robustness of the various ensemble methods compared to using just one model. Given our ensemble detectors, we consider the following options: (a) PPL detection for each model; (b) (Fast)DetectGPT for each model; (c) Binoculars score for each pair of models; (d) PPL detection with average PPL scores or with the log-probabilities of the optimal distribution $q^{*}$; [^8] Strictly speaking, not a perplexity, as the Blahut-Arimoto weights change for each token. (e) MOSAIC; (f) the MOSAIC score, using a uniform weighting scheme instead of Blahut-Arimoto weights. Methods (a,b) require just one model; (c) requires two; (d,e,f) require four. For the sake of space, we only report in Table 3 for methods (a) and (b) the model with the best performance on the CC_News subset of Binoculars; likewise, for (c) we select the best pair of models on the same test corpus (TowerBase-7b as the detector, and Llama-2-7b as the auxiliary model). The complete set of results is in Table 5 in Appendix E.
We see a large variation across datasets and generation techniques: for some, a near-perfect detection can be achieved (notably Binoculars datasets, and to a lesser extent, Ghostbuster, for which PPL-based detectors can be remarkably good). However, varying the domain (Scholarly texts) and/or languages (M4) can have a detrimental effect on detection performance.
Among the baselines (Table 2), Binoculars is the most robust and achieves the best average performance, perhaps owing to the better underlying detector model (Falcon vs. GPT2). Binoculars is very sensible to changes in scripts, as the results on the M4 dataset for Arabic (Arabic script), Bulgarian and Russian (both written in Cyrillic), show.
Our best single-model detector (FastDetectGPT with Tower-13b, selected out of 12 alternatives) achieves very good average scores and outperforms the original Binoculars baseline. By contrast, our best implementation of the Binoculars score with 2 models from our ensemble (also selected out of 12 combinations), is less effective and lags behind the Falcon-based Binoculars detector. Using 4 models, MOSAIC(avg) yields the best on average, and dispenses with a prior search for the optimal configuration. For RAID, arguably the more challenging dataset, it compares with the best Binoculars model from Hans et al. (2024).
An interesting follow-up question is about the respective strength of each detector model: can this be measured using the Blahut-Arimoto weights? For our ensemble, we find that all models get a reasonable (and varying) share of these weights. We see however that Llama-7b instruct consistently receives very large weights (see Figure 4(a) in Appendix). For the multilingual tests from M4, we observe that the weights of the TowerBase models tend to increase, as compared to when just looking at English texts (see Figure 2 in Appendix). This illustrates the benefits of using complementary models, each one with its own domain of “expertise”.
@@ 4.1 The robustness of ensemble methods
Scholarly texts pose another problem. As these were produced by a GPT-2 model fine-tuned on Arxiv’s “computation and language” section, the distribution of tokens greatly differs from all models considered in this study. The generated texts therefore appear completely out-of-domain and get a high perplexity for all models, which makes them difficult to discriminate from human texts (see Figure 3 in the Appendix) with perplexity-based models. This holds similarly for our ensemble (e.g., MOSAIC-4) and for baseline models (e.g., Falcon-based Binoculars), which achieve respective AUROCs of 0.421 and 0.505. This shows that brittleness issues are not fully solved, and hints at including more domain-adapted models in detector ensembles.
Adversarial attacks, reported in Table 7, deteriorate slightly the performances of our method with the exception of "synonym replacement", which makes generated texts more surprising while not influencing human-produced texts in the same way (the idea behind DetectGPT), completely breaking our method. Interestingly, replacing some characters with homoglyphs actually improves our results, we hypothesize it is because the Tower models have seen Cyrillic in their training data.
4.2 Including the generator in the detector ensemble
Table 4 reports results with Binoculars+, where we augment Binoculars with comparable texts generated with two models: Llama-7b, which is part of our ensemble, and Mistral-7b, which is not. For Pubmed and CC_news, we see the same trend: both the original artificial texts and the Mistral-generated texts are much easier to detect than Llama-generated texts. Contrarily, for CNN, Llama-7b is almost perfectly detected. Overall, having the generator inside the ensemble of detectors does not seem particularly advantageous.
However, it should be noted that the sampling method plays an important role in the performance of unsupervised detection methods, as displayed in Table 6 in the Appendix. Recall that our regeneration uses a temperature of $0.6$ along with nucleus sampling ($p=0.9$) and a repetition penalty of 1.18 (see Section 3.1), as recommended by the official Llama repository [^9] https://github.com/meta-llama/llama/blob/main/example_text_completion.py as they produce more diverse, and in a way, more human-like texts than when using greedy decoding. This means that the actual sampling distribution does not fully match the base distribution of the detector, and consistently generates ”surprising” tokens for the generator model.
In a follow-up experiment, we replace the distribution computed by Llama-2-7b with a “distorted” version, which approximates the effects of temperature and nucleus-sampling [^10] We divide the model logits by the temperature, then apply softmax and perform nucleus filtering with $p=0.9$. The resulting distribution is smoothed so out-of-nucleus tokens still get a small positive value.. This has a clear effect on the Blahut–Arimoto weights defining $q^{*}$, which increase for Llama-2-7b and decrease for Llama-2-7b-instruct, but hardly change the detection performance (see Figure 4 in Appendix).
Table 4: Identification results for the Binoculars and Binoculars+ datasets. Single models detectors are PPL-based. AUROC scores. Best scores in each column are in boldface.
Pubmed CNN CC news Avg.
Orig. Llama Mist. Orig. Llama Mist. Orig. Llama Mist.
TowerBase-7b 0.798 0.631 0.806 0.557 0.999 0.639 0.582 0.573 0.669 0.695
TowerBase-13b 0.823 0.528 0.801 0.561 0.999 0.641 0.578 0.439 0.652 0.669
Llama-2-7b 0.786 0.676 0.820 0.544 1.000 0.704 0.556 0.617 0.693 0.711
Llama-2-7b-chat 0.817 0.690 0.843 0.598 1.000 0.716 0.617 0.632 0.718 0.737
$q^{\star}$ (log-probs) 0.807 0.641 0.836 0.566 1.000 0.727 0.579 0.568 0.703 0.714
MOSAIC-4 (avg) 0.992 0.887 0.961 0.993 0.999 0.971 0.971 0.854 0.940 0.952
4.3 Augmenting the ensemble
With a Strong Model.
MOSAIC makes the augmentation of the ensemble quite easy, as long as all models use the same underlying tokenizer. To showcase the effect of this feature, we add Phi-3 (Abdin et al., 2024) in our ensemble and experiment with Binoculars and Ghostbuster datasets. In this experimental setting, this extension of the ensemble is of little consequence for the former test, and provides clear gains for the latter (see lines “+phi” in Appendix, Table 9). Accordingly, we observe that Phi consistently gets a substantial share of the Blahut–Arimoto weights (see Figure 5 in Appendix) and plays a significant role in the classification decision.
With a Weak Model.
As noted by Hans et al. (2024), simple random generators are hard to detect for Binoculars (as well as for detectors using a PPL-based threshold): this is because random word salads are “surprising” (have a high perplexity) for well-trained detectors, and tend to be confused with human productions. We reproduce this observation using the corpus generated with a unigram model and observe that all baselines, as well as our proposed detector, achieve AUROC scores close to 00 for this dataset. The PPL of such “word salads" is however much larger than for human texts, so setting an upper bound of the human PPL would, in that case, provide a very easy fix.
It is tempting to see whether adding such a poor generator into our combination would be of any help. For this experiment, we reuse the generator unigram language model and combine it with our four baseline detectors. The detector obtained with this extended ensemble remains unable to sort random from human texts: having the random model in the MOSAIC algorithm does not make random texts more likely. The added unigram model is also slightly detrimental for detecting strong generators, as we observe a mild drop in performance compared to using just 4 models (see Appendix, Table 9). This is because the unigram model predicts unexpected tokens most of the time and therefore often gets a significant weight in the Blahut–Arimoto weights (see Figure 6 in Appendix), which leads to a huge discrepancy between the cross-entropies scores that are averaged in Definition 1. However, this effect remains small; even with this weak model added, our ensemble detector remains rather strong.
@@ 5 Related Work
The improved text generation abilities of LLMs raise concerns about potential misuses such as disinformation (Zellers et al., 2019), abusive content (Crothers et al., 2023), forged academic publications (Liu et al., 2024), or cheating during exams (Vasilatos et al., 2023). Since such fake texts seem difficult for humans to spot (Gehrmann et al., 2019), the issue of automatically detecting machine-generated texts has been subject to an increasing focus. This problem can be framed as a binary human vs. non-human decision, as the problem of detecting one known artificial agent (e.g.,  ChatGPT (Mitrović et al., 2023; Liu et al., 2024)), or as discriminating the correct model in a predefined list (Li et al., 2023). Another distinction is between closed-domain (e.g. scientific (Liyanage et al., 2022), academic (Liu et al., 2024) or user-generated content (Fagni et al., 2021; Kumarage et al., 2023)) vs. open-domain text detection. Assuming the generator models are known, various settings can be considered, depending on whether models can be openly queried (open parameters), whether they expose their full logits, or just the top prediction (and associated probability), etc.
@@
Supervised detection with a single generator often achieves detection rates in the high 90s (Zellers et al., 2019; Guo et al., 2023; Liu et al., 2024), using classifiers based on Roberta (Conneau et al., 2020) or T5 (Raffel et al., 2020). However, these approaches are brittle and their success depends on particular generator-detector pairs (Antoun et al., 2024), prompting e.g. Verma et al. (2023) to design automatic feature extractors from multiple detectors to improve the robustness of their system.
@@
Unsupervised detection is more challenging. Most approaches rest on the idea that human-written texts are more “surprising” than artificial texts [^11] Assuming generation does not use random sampling, in which case the reverse is likely to be observed, as long artificial texts drift away from natural writings (Zellers et al., 2019)., leading to a difference in token-wise log-probability121212(Mitchell et al., 2023) argues that the difference is better seen at the level of log-ranks.. This idea is used in GPTzero [^13] https://gptzero.me/ and thresholding perplexity usually provides strong baselines (see, inter alia, (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023)). Such techniques heavily rely on the detector model(s) used to compute the log-probabilities of input texts, which must be robust to variations in domains, genres, styles, and languages (Wang et al., 2024a); and to variations in the generator itself (Antoun et al., 2024).
@@
Mitchell et al. (2023) and Bao et al. (2024) exploit a similar intuition, arguing that small random perturbations of an artificial text will on average make it less likely, unlike human-written texts. They derive a statistical criterion based on the curvature of the log-probability function, and report near-perfect detection scores on three types of texts, generated by 5 models. The Binoculars score of Hans et al. (2024) also relies on a function of the per-token log-perplexity, contrasted with the cross-entropy of an auxiliary model.
@@
Abandoning generator-detector-based techniques altogether, (Mao et al., 2024; Yang et al., 2024) develop effective detection approaches based on regeneration, prompting the (known) generator to regenerate part of the input text. The intuition is that artificial inputs are likely to be regenerated exactly, while human texts exhibit greater redundancy, resulting in the need for longer codelengths. Other strategies include text watermarking (Kirchenbauer et al., 2023a, b; Liu and Bu, 2024), though its efficiency and robustness are still subject to discussions, e.g., (Zhang et al., 2023).
Recent works focus on detection robustness. Wang et al. (2024b) find that after simple modifications, only watermarking remains able to accurately identify artificial documents. Dugan et al. (2024) present artificial texts generated with multiple models and sampling strategies, additionally subject to various adversarial attacks, observing that most detectors suffer large drops in performance. In their comparison, Binoculars (Hans et al., 2024) stands out, achieving decent detection scores at False Positive Rates under 1%.